OpenStudy (anonymous):

use rational exponents and simplify ³√x^15

OpenStudy (callisto):

\[\sqrt[3]{x^{15}}={(x^{\frac{1}{3}})^{15}} = {x^{\frac{1}{3}\times 15}} =?\]

OpenStudy (anonymous):

x 46/3 ?

OpenStudy (callisto):

Nope...

OpenStudy (callisto):

what is the value of (1/3) x 15?

OpenStudy (anonymous):

15/3 I didn't get this in high school and this is ten times crazier

OpenStudy (anonymous):

reduce it down to 5 ????

OpenStudy (callisto):

Yes!~

OpenStudy (callisto):

So, it becomes x^5 and that's your answer!~